OpenStudy (anonymous):

Given the function f(x) = The quantity of 3x minus 4, divided by 5, which of the below expressions is correct? f−1(x) = The quantity of 5x plus 4, divided by 3. f−1(x) = The quantity of 5x minus 4, divided by 3. f−1(x) = The quantity of negative 3x minus 4, divided by 5. f−1(x) = The quantity of 4 minus 3x, divided by 5.

OpenStudy (anonymous):

@AJ01

OpenStudy (anonymous):

Is this the function, \(\large f(x)=\frac{3x-4}{5}\)?

OpenStudy (anonymous):

So we have \(\large y=\frac{3x-4}{5}\). To find the inverse function, switch x and y. So now we have \(\large x=\frac{3y-4}{5}\) <=> \(5x+4=3y\) <=> \(\large y=\frac{5x+4}{3}\) So the inverse function will be \(\large f^{-1}(x)=\frac{5x+4}{3}\). So A is correct